OpenStudy (anonymous):

36% of what number is 18?

OpenStudy (kc_kennylau):

Let that number be \(x\).

OpenStudy (kc_kennylau):

\(x\times36\%=18\)

OpenStudy (kc_kennylau):

can you continue? :)

OpenStudy (anonymous):

6.48

OpenStudy (anonymous):

write the equation and solve

OpenStudy (anonymous):

change 36% into a decimal, 0.36 then set up the equation: .36*x=18, then solve for x by dividing both sides by .36

OpenStudy (anonymous):

um... it has to be greater than 18 @musa19965

OpenStudy (jdoe0001):

\( \begin{array}{llll} quantity&percentage\\ x&100\\ 18&36 \end{array}\implies \cfrac{x}{18}=\cfrac{100}{36}\)

OpenStudy (jdoe0001):

lol

OpenStudy (anonymous):

ooops Solve for x: 0.36 x = 18 Divide both sides by a constant to simplify the equation. Divide both sides of 0.36 x = 18 by 0.36: (0.36 x)/0.36 = 18/0.36 Any nonzero number divided by itself is one. 0.36/0.36 = 1: x = 18/0.36 Express 18/0.36 in decimal form. 18/0.36 = 50.: Answer: | | x = 50.

OpenStudy (anonymous):

I got it :) thanks everyone

OpenStudy (anonymous):

not supposed to do it for them @musa19965

OpenStudy (anonymous):

ooops sorry